Marsha Jones has bought a used Mercedes horse transporter for her Connecticut estate. It cost $50,000. The object is to save on horse transporter rentals. Mars a had been renting a transporter every oter ee r S2 15 per week plus $1.75 per mile. Most o t erips are o miles n otal Mar na su ives ed era 54 ·With the ne tr spor er she wil have o a for diesel fuel and maintenance, at about $.60 per mle. Insurance costs for Marshas transporter are $1,950 per year. The trans orter wi robab be worth 30 000 in real terms arter eart ears, when Mars as horse Nike wi ї be ea t retire. Ass me a no na discount rate 017% and a recast in ation rate of 3 Marshas rans orter is a personal outlay, not a business or financial investment, so taxes can be ignored. Hint: All numbers given in the questions are in real terms. Assume CF at end of year, for simplicity. Calculate the NPV of the investment. (Do not round intermediate calculations. Round your answer to the nearest whole dollar amount.) NPV
Answer #1

Yr0 CF0 = (50,000)

This is cash outlay for new transporter

Yr1 CF1 will be incremental cash flows. So we will find cost as per original method and then cost as per new transporter.

Original cost per week = 215 + (1.75*90) + 40 = 412.5

Original cost p.a. = 21450

New transporter cost = 0.6*90*52 + 1950 = 4758

So we have cash saving of = 21450 - 4758 = 16692 for all 8 yrs.

Scrap value in yr 8 = 30000

So cash flows are :

Cf0 = (50000)

Cf01 = 16692

F01 = 7

Cf02 = 46692

I = 7

Npv = 67133